Re-writing the following equation in the form y=(x-h)^2+k using "complete the square" and find the vertex and intercepts.
y=x^2+3x+5

Answer by lwsshak3(11628)

Re-writing the following equation in the form y=(x-h)^2+k using "complete the square" and find the vertex and intercepts.
y=x^2+3x+5
y=(x^2+3x+9/4)-9/4+5
y=(x+3/2)^2+11/4
..
vertex: (-3/2,11/4)
y-intercept
ser x=0
f(0)=5
y-intercept: (0,5)
..
x-intercept: none (range: [5, ∞)